# bibliography.bib

### v1

@@ I Introduction
To run large-scale quantum algorithms [dalzell2023quantum], it is essential to incorporate quantum error correction and fault tolerance to ensure reliable results [Shor1995, Steane1996, shor1996fault, kitaev1997quantum, preskill1998reliable]. In a fault-tolerant setting, logical qubits are encoded using a quantum error-correcting code and logical operations are designed to limit the propagation of errors. Consequently, executing quantum algorithms on any quantum computing platform requires translating logical circuits into physical ones compatible with hardware-specific constraints—a process known as compilation.
Most state-of-the-art research in fault-tolerant compilation [herr_lattice_2017, zhu_ecmas_2023, lao_mapping_2018, watkins_high_2024, silva_multi-qubit_2024, beverland_surface_2022, molavi_dependency-aware_2024] has focused on the surface code architecture [Kitaev_2003, Dennis_2002, fowler_surface_2012], which can be implemented with a planar layout of qubits with nearest-neighbor connectivity. A key challenge in this setting is the implementation of logical two-qubit gates, which may naively require non-local qubit connectivity. A standard technique to address this issue is lattice surgery [horsman_surface_2012, landahl_quantum_2014, litinski_game_2019], where logical qubits are made to interact by performing measurements of additional ancilla qubits placed in the region connecting the surface code patches that encode these logical qubits.
Alternatives to the surface code can be found within the broader class of topological codes, which includes the color code [bombin_topological_2006, kubica_abc_2018] and the folded surface code [kubica_unfolding_2015, moussa_transversal_2016]. These quantum error-correcting codes can also be implemented using a planar layout of qubits with geometrically-local (rather than strictly nearest-neighbor) connectivity. In return, the qubit overhead associated with performing error correction may be substantially reduced. Despite these potential gains, fault-tolerant compilation for general topological codes remained largely underexplored.
In this work, we address the problem of fault-tolerant compilation for topological codes by introducing a general formalism that is applicable beyond the surface code and spans multiple levels of abstraction as illustrated in Fig. 1. On the microscopic level, we define the concept of a code substrate—a blueprint for realizing a topological code and its corresponding lattice surgery scheme; see Fig. 1b. We present two concrete examples of such substrates that support the color code and the folded surface code. While the folded surface code requires a substrate with slightly more complex qubit connectivity than the color code, it is expected to achieve higher fault-tolerant thresholds while retaining key advantages of the color code, such as transversal single-qubit Clifford gates.
@@ II Main Ideas and Problem formulation
In this work, we consider the problem of compiling logical circuits consisting of Clifford$+\mathrm{T}$ gates to lattice surgery operations assuming that $\mathrm{T}$ gates are realized using magic state distillation and injection [bravyi_universal_2005, lee_low-overhead_2024]. However, we do not assume any specific distillation procedure.
Previous work on fault-tolerant computation such as [beverland_surface_2022, zhu_ecmas_2023, molavi_dependency-aware_2024, lao_mapping_2018, watkins_high_2024, silva_multi-qubit_2024], focused on the surface code and assumed a rectangular logical qubit connectivity. To consider the compilation problem more generally, i.e., without assuming any particular qubit connectivity and quantum error-correcting code beforehand, we distinguish the microscopic and macroscopic levels of abstraction. At the microscopic level, the main problem is how to realize a lattice surgery scheme that implements logical two-qubit operations. At the macroscopic level, a central task is to assign logical qubits to physical locations, obtaining a logical routing graph whose shape is dictated by concrete choices at the microscopic level. Then valid paths between logical data qubits have to be identified to implement the logical quantum circuit.
To showcase the resulting formalism, consider the example of compilation with the 2D color code illustrated in Fig. 1. We start from the assumption of some geometrically local connectivity of physical data qubits (black circles) and ancilla qubits (white circles) as illustrated by the graph in Fig. 1a. In particular, we assume a stacked planar architecture that consists of a constant number of planar layers with qubit connectivity that is local within each layer and between layers (but not necessarily nearest-neighbor). Depending on the hardware it may be more favorable to either consider the stacked layers as is or project them onto one layer while allowing for a more complex (albeit still local) qubit connectivity.
@@
Building on this, we split the substrate into several regions (black triangles in Fig. 1b) that are dedicated to support logical data qubits, magic state patches, and the remaining routing space. For concreteness, we assume that the magic states are encoded in the same code as the logical qubits. Moreover, it is instructive to think of the routing space as being further subdivided into several ancilla patches. A logical data patch $L_{i}$ and three ancilla patches are illustrated in Fig. 1b. The ancilla regions form the building blocks of what we refer to as snakes on the substrate, which are ancilla regions that are used to implement lattice surgery schemes between logical data patches. Note that a logical patch may also correspond to a logical ancilla qubit placed in the routing space.
A concrete assignment of logical data patches, magic state patches, and ancilla regions on the substrate determines a logical routing graph as shown in Fig. 1c, whose vertices correspond to logical data patches, magic state patches, and ancilla patches of the routing space. The edges of the routing graph indicate the connectivity between the regions on the substrate. Paths on this macroscopic routing graph correspond to microscopically defined snakes together with a logical ancilla patch that are used to perform $\mathrm{CNOT}$ gates between logical data patches.
@@ III-A Topological Quantum Codes
Topological codes [bombin_topological_2006, bombin_introduction_2013, bombin_topological_2010] are a class of quantum error-correcting codes that are defined by placing qubits on some manifold; we restrict our attention to stabilizer codes [gottesman1998heisenberg]. In particular, in two dimensions checks and data qubits are often associated with faces and vertices of some tiling. Each check of the stabilizer group $S$ is local, supported only on a constant number of qubits, whereas logical qubits are encoded non-locally. Prominent examples for topological codes are color codes and surface codes that are defined using different tilings of 2D surfaces.
Two-dimensional color codes [bombin_topological_2006] are defined on 3-colorable and 3-valent tilings. An important family of color codes are triangular color codes on a hexagonal tiling whose faces and boundaries are three-colored. An example is displayed in 2(a). Here, each face contained within the triangular region defines both an $X$ and a $Z$ check. $Z$ and $X$ logical Pauli string operators [bombin_topological_2006] terminate at boundaries of the same color as displayed in the figure, or connect three boundaries of different colors. The code distance $d$ corresponds to the minimum weight of a non-trivial logical operator. In the example in 2(a) the distance is $d=5$.
The surface code [Kitaev_2003, Dennis_2002, fowler_surface_2012] can be defined on a rectangular tiling with two different types of boundaries. The checks are placed in a checkerboard pattern. An example is shown in 2(b). Here, representatives of logical operators can be defined as strings that connect boundaries of the same type. In this example, horizontal strings correspond to $Z_{L}$ (connecting so-called rough boundaries) and vertical strings (connecting so-called smooth boundaries) to $X_{L}$. Therefore, the distance of the displayed example is $d=5$. We note that a rotated variant of the surface code, which is supported within a region enclosed by the dotted line in 2(b), leads to a smaller qubit overhead.
In its original formulation, the surface code does not admit transversal implementation of the logical $\mathrm{S}$ gates that are compatible with geometrically-local qubit connectivity. This issue can be circumvented by folding [moussa_transversal_2016, kubica_unfolding_2015] the surface code along the diagonal (orange nodes in 2(b) and 2(c)) such that faces of different check type are stacked on top of each other. Additionally, smooth and rough boundaries are also stacked, such that each boundary enables access to both logical $X_{L}$ and $Z_{L}$ representatives, similarly as for the color code.
@@ III-B Lattice Surgery
Quantum computation requires a universal gate set, such as $\{\mathrm{CNOT},\mathrm{H},\mathrm{T}\}$. In general, transversal implementation of the two-qubit $\mathrm{CNOT}{}$ gate requires a non-local qubit connectivity. Therefore, to implement $\mathrm{CNOT}{}$ gates with local connectivity, one may resort to lattice surgery [horsman_surface_2012].
@@ III-C Related Work
Let us conclude this section by reviewing how the resulting macroscopic compilation task is solved in related work. Surface code compilation with a universal or smaller gate set was explored in [herr_lattice_2017, zhu_ecmas_2023, lao_mapping_2018]. Furthermore, it has been shown that the complexity of optimizing a quantum circuit in the lattice surgery model is NP-hard [herr_optimization_2017]. Related work by Litinski [litinski_game_2019] motivated further research [watkins_high_2024, silva_multi-qubit_2024], though it was argued that commuting Clifford gates until the end of the circuit may limit the parallelism [beverland_surface_2022, molavi_dependency-aware_2024]. The aforementioned works focus on surface codes only; consequently, fault-tolerant compilation for other codes remained underexplored, despite some early ideas [litinski_combining_2017].
Using color codes instead of surface codes offers several advantages. For instance, the color code has a higher qubit density, supports transversal Clifford gates and hosts both $X_{L}$ and $Z_{L}$ per boundary, which simplifies the routing to some extent by removing the constraint of connecting a specific choice of boundaries. However, decoding algorithms for the color code tend to perform worse [Kubica2023, chamberland_triangular_2020, sahay2022decoder, gidney2023new, lee_color_2025]. Nevertheless, recent experiments with color codes [postler2022demonstration, ryan2022implementing, rodriguez2024experimental, pogorelov2025experimental] and importantly, lattice surgery demonstration [lacroix_scaling_2024] underpin the potential of color codes as a viable fault-tolerant architecture.
Concerning the surface code, permitting a stacked planar architecture or relaxing the constraint on nearest-neighbor qubit connectivity enables to use the folded surface code [kubica_unfolding_2015, moussa_transversal_2016]. It retains many advantages of the color code while allowing straightforward decoding, the same way as for the standard surface code [Dennis_2002, higgott_sparse_2023]. However, to the best of our knowledge, no previous work on compilation with these codes exists.
@@ IV Formalism for Fault-Tolerant Compilation
On the physical level we assume that the local qubit connectivity is given by a graph (cf. Fig. 1a) that may have a constant number of layers of physical qubits such that qubits within each layer and between neighbouring layers can have geometrically local interactions. Note that such an architecture is less hardware-demanding than the one needed to implement bivariate bicycle codes [bravyi_high-threshold_2024, kovalev2013quantum]; as such, it may be easier to realize with superconducting qubits [Devoret2013, Blais2021], as well as neutral atoms [Saffman2010, Browaeys2020] or trapped ions [Cirac1995, Leibfried2003]. Physical data and ancilla qubits (black and white in Fig. 1) are placed on the vertices and their connectivity is given by the edges. This is the basis for choosing the code substrate, which is a tiling $\mathcal{S}$ (Fig. 1b) that can be used to realize logical qubits encoded in some topological quantum code. The substrate allows us to view a concrete physical qubit connectivity abstractly. We use the substrate to illustrate the layout of data qubits and geometrically-local checks of the code (faces of $\mathcal{S}$).
@@
In this work we assume that the input circuit $\mathcal{C}$ is a logical quantum circuit that contains logical $\mathrm{CNOT}+\mathrm{T}$ gates. To obtain a universal gate set, single-qubit Clifford gates would be necessary as well, but they are not considered here since the computational hardness of the routing challenge for topological codes is dominated by multi-qubit operations [beverland_surface_2022, molavi_dependency-aware_2024, silva_multi-qubit_2024, zhu_ecmas_2023]. Moreover, our concrete choices of codes and substrates detailed below allow for transversal single-qubit Clifford operations.
To realize logical two-qubit operations we use lattice surgery schemes between logical patches on the substrate. Upon a choice of $\mathcal{L},\mathcal{A},\mathcal{F}$ and a code, lattice surgery requires the implementation of specific stabilizer measurements to connect logical patches as reviewed in Sec. III-B. To this end we realize what we refer to as snakes on the substrate. Snakes consist of neighboring ancilla regions and potentially additional elements from $\mathcal{A}\backslash\mathcal{A}^{*}$ and allow for the merge of distant logical (data/ancilla) patches to measure joint two-qubit logical Pauli operators. Throughout we assume that paths on $\mathcal{R}$ between distant logical data qubits contain both a snake as well as a logical ancilla patch, together with potentially needed physical ancilla qubits from $\mathcal{A}\backslash\mathcal{A}^{*}$ to perform a logical $\mathrm{CNOT}$ gate.
The concrete design of the stabilizers of a snake requires special care as the protocol must be implemented in a fault-tolerant and distance-preserving way. We detail two concrete constructions for logical color code patches and color code snakes, and (folded) surface code patches and surface code snakes in Sec. VI-B and Sec. VI-C, respectively.
@@ V Macroscopic Compilation of $\mathrm{CNOT}+\mathrm{T}$ circuits
Figure 5: Considered layouts, i.e., a choice of $\mathcal{L}$ and $\mathcal{A}$ that determines the routing graph $\mathcal{R}$, and their asymptotic packing ratio $c$, which we define as the number of logical data qubit patches (assuming that the folded surface code forms one patch) to the total number of patches. Light blue and gray nodes depict logical data qubits and ancilla patches, respectively. Due to substrate constraints, routes between neighboring logical data patches with one edge only are not allowed.
@@
In the following sections we detail how we can solve the routing and mapping problem for the macroscopic compilation.
@@ V-A Shortest-First VDP Solving
Let us start by considering logical circuits with $\mathrm{CNOT}{}$ gates between disjoint pairs of qubits only. Given a routing graph $\mathcal{R}=(V_{\mathcal{R}},E_{\mathcal{R}})$ and a labeling of logical data patches $\Lambda^{*}$, one can formalize the search for paths as vertex disjoint path problem (VDP) [chuzhoy_approximating_2015, chuzhoy_improved_2016, bacco_shortest_2014]. The aim of the VDP is to find non-overlapping paths between pairs of vertices $C=\{(r_{1},r^{\prime}_{1}),...,(r_{k},r^{\prime}_{k})\mid r_{i}\neq r_{j},r^{\prime}_{i}\neq r^{\prime}_{j},r_{i}\neq r^{\prime}_{j}\,\forall i,j\}\subset V_{\mathcal{L}}\times V_{\mathcal{L}}$. The considered vertices in $V_{\mathcal{L}}$ are mapped to logical qubit labels via a choice of $\Lambda^{*}$. Hence, the pairs to connect—i.e., the elements of $C$—represent the control and target qubits of $\mathrm{CNOT}{}$ gates and the paths contain the snakes and logical ancillas to perform lattice surgery between pairs of logical data qubits. Overall, $C$ can be considered as quantum circuit consisting of $\mathrm{CNOT}{}$ gates.
@@
A simple greedy algorithm [chuzhoy_approximating_2015] to find solutions for the VDP is to start with an empty solution $W_{i}$ that is iteratively filled with paths $p_{j}$. We start by computing the shortest path with Dijkstra’s algorithm between each pair in $C$, choose the shortest among those paths and add it to the solution $W_{i}=W_{i}\cup\text{argmin}_{p_{i}\in\{\textsc{DIJKSTRA}(r_{j},r^{\prime}_{j})\mid(r_{j},r^{\prime}_{j})\in C\}}|p_{i}|$. Then the vertices in the chosen path $p_{i}$ are removed from $\mathcal{R}$. As pointed out in the beginning of the section, a substrate may require nontrivial constraints on paths on $\mathcal{R}$ to be valid such that Dijkstra may have to be adapted accordingly. This process of finding the shortest path and removing vertices is repeated until paths for all pairs in $C$ are found or no further non-overlapping path can be found. It is guaranteed that no overlapping paths can occur since consumed vertices are removed from $\mathcal{R}$. Running this shortest-first VDP subroutine yields a set of paths $W_{i}=\{p_{1},...p_{n}\}$ that correspond to logical two-qubit gates, which can be executed in parallel with lattice surgery on the substrate.
@@
In general, a $\mathrm{CNOT}{}$ quantum circuit contains multiple sets of pairs $C_{i}$ as we do not assume that all $\mathrm{CNOT}{}$ gates have disjoint qubit support. Thus a $\mathrm{CNOT}{}$ quantum circuit can be viewed as $\tilde{C}=\{C_{1},...,C_{m}\}$. Additionally, the routing graph $\mathcal{R}$ may not allow us to find a complete set of vertex-disjoint paths for a given $C_{i}$. Thus, the aforementioned subroutine must be embedded in an algorithm, which we call shortest-first VDP solving. We start by employing the shortest-first VDP subroutine for $C_{1}$ and add the resulting set of paths $W_{i}$ to the overall solution $C^{\prime}$. If not all pairs in $C_{1}$ can be connected by non-overlapping paths, the remaining elements from $C_{1}$ are copied to $C_{2}$. This may lead to $C_{2}$ containing gates with shared qubit labels. Thus, $C_{2}$ must be fixed by pushing problematic gates into $C_{3}$, such that $C_{2}$ contains gates with disjoint qubit labels only. Then, $C_{3}$ and the subsequent sets may be adapted in the same manner. Note that pushing requires us to respect the order of gates, as otherwise we could shuffle non-commuting gates. This shortest-first VDP subroutine together with the pushing is repeated until a path for each pair is found. The solution contains multiple sets of paths $C^{\prime}=\{W_{1},...W_{\Delta}\}$, where $\Delta$ is the depth of the final, compiled circuit.
@@
To include $\mathrm{T}$ gates in this picture, one can expand the shortest-first VDP subroutine to allow singleton tuples, i.e., single vertices $r_{\ell}\in V_{\mathcal{L}}$ corresponding to logical qubit labels via $\Lambda^{*}$. Additionally, the vertices of the magic state patches $V_{\mathcal{F}}$ must be taken into account in the routing problem formulation. We model magic state patches as logical patches that “receive” magic states after some fixed reset time from some adjacent magic state factory. The $\mathrm{T}$ gate can be implemented by gate injection from any available magic state patch. Thus, one can expand the shortest-first VDP subroutine by greedily computing all paths from the logical data qubit on which a $\mathrm{T}$ gate should be applied, $r_{\ell}$, to the magic state patches from $V_{\mathcal{F}}$, if their $\mathrm{T}$ state is available. The shortest paths between $r_{\ell}$ and the available magic state patches is compared to all other paths corresponding to elements in $C_{i}$ and the shortest is added to the solution as before. The availability of magic states is modeled by defining a reset time $t\in\mathbb{Z}$, which is initialized as positive integer and reduced by $1$ after each completion of the shortest-first VDP subroutine. The factory is available when $t\leq 0$ and reset to the original value if the magic state was consumed. Note that is a pessimistic assumption as the units of time required to create a magic state may be less coarse grained.

### v2

# Lattice Surgery Compilation

@@ I Introduction
To run large-scale quantum algorithms [1], it is essential to incorporate quantum error correction and fault tolerance to ensure reliable results [2, 3, 4, 5, 6]. In a fault-tolerant setting, logical qubits are encoded using a quantum error-correcting code and logical operations are designed to limit the propagation of errors. Consequently, executing quantum algorithms on any quantum computing platform requires translating logical circuits into physical ones compatible with hardware-specific constraints—a process known as compilation.
Most state-of-the-art research in fault-tolerant compilation [7, 8, 9, 10, 11, 12, 13] has focused on the surface code architecture [14, 15, 16], which can be implemented with a planar layout of qubits with nearest-neighbor connectivity. A key challenge in this setting is the implementation of logical two-qubit gates, which may naively require non-local qubit connectivity. A standard technique to address this issue is lattice surgery [17, 18, 19], where logical qubits are made to interact by performing measurements of additional ancilla qubits placed in the region connecting the surface code patches that encode these logical qubits.
Alternatives to the surface code can be found within the broader class of topological codes, which includes the color code [20, 21] and the folded surface code [22, 23]. These quantum error-correcting codes can also be implemented using a planar layout of qubits with geometrically-local (rather than strictly nearest-neighbor) connectivity. In return, the qubit overhead associated with performing error correction may be substantially reduced. Despite these potential gains, fault-tolerant compilation for general topological codes remained largely underexplored.
In this work, we address the problem of fault-tolerant compilation for topological codes by introducing a general formalism that is applicable beyond the surface code. This spans multiple levels of abstraction as illustrated in Fig. 1. On the microscopic level, we define the concept of a code substrate—a blueprint for realizing a topological code and its corresponding lattice surgery scheme; see Fig. 1b. We present two concrete examples of such substrates that support the color code and the folded surface code. While the folded surface code requires a substrate with slightly more complex qubit connectivity than the color code, it is expected to achieve higher fault-tolerant thresholds while retaining key advantages of the color code, such as transversal single-qubit Clifford gates.
@@ II Main Ideas and Problem formulation
[…]
Previous work on fault-tolerant computation such as [12, 8, 13, 9, 10, 11], focused on the surface code and assumed a rectangular logical qubit connectivity. To consider the compilation problem more generally, i.e., without assuming any particular qubit connectivity and quantum error-correcting code beforehand, we distinguish the microscopic and macroscopic levels of abstraction. At the microscopic level, the main problem is how to realize a lattice surgery scheme that implements logical two-qubit operations. At the macroscopic level, the central task is to assign logical qubits to physical locations, obtaining a logical routing graph whose shape is dictated by concrete choices at the microscopic level. Then valid paths between logical data qubits have to be identified to implement the logical quantum circuit.
To showcase the resulting formalism, consider the example of compilation with the 2D color code illustrated in Fig. 1. We start from the assumption of some geometrically local connectivity of physical data qubits (black circles) and ancilla qubits (white circles) as illustrated by the graph in Fig. 1a. In particular, we assume a stacked planar architecture that consists of a constant number of planar layers with qubit connectivity that is local within each layer and between layers (but not necessarily nearest-neighbor). Depending on the hardware it may be favorable to use indeed a stacked planar architecture or to project the stacked layers into one while accepting a more complex qubit connectivity.
@@
Building on this, we split the substrate into several regions (black triangles in Fig. 1b) that are dedicated to support logical data qubits, magic state patches, and the remaining routing space. For concreteness, we assume that the magic states are encoded in the same code as the logical qubits. Moreover, it is instructive to think of the routing space as being further subdivided into several ancilla patches. A logical data patch $L_{i}$ and three ancilla patches are illustrated in Fig. 1b. The ancilla regions form the building blocks of what we refer to as snakes on the substrate. They are ancilla regions that are used to implement lattice surgery schemes between logical data patches. Note that a logical patch may also correspond to a logical ancilla qubit placed in the routing space.
A concrete assignment of logical data patches, magic state patches, and ancilla regions on the substrate determines a logical routing graph as shown in Fig. 1c. Its vertices correspond to logical data patches, magic state patches, and ancilla patches of the routing space. The edges of the routing graph indicate the connectivity between the regions on the substrate. Paths on this macroscopic routing graph correspond to microscopically defined snakes together with a logical ancilla patch that are used to perform $\mathrm{CNOT}$ gates between logical data patches.
@@ III-A Topological Quantum Codes
Topological codes [20, 26, 27] are a class of quantum error-correcting codes that are defined by placing qubits on some manifold; we restrict our attention to stabilizer codes [28]. In particular, in two dimensions checks and data qubits are often associated with faces and vertices of some tiling. Each check of the stabilizer group $S$ is local, supported only on a constant number of qubits, whereas logical qubits are encoded non-locally. Prominent examples for topological codes are color codes and surface codes that are defined using different tilings of 2D surfaces.
Two-dimensional color codes [20] are defined on 3-colorable and 3-valent tilings. An important family of color codes are triangular color codes on a hexagonal tiling whose faces and boundaries are three-colored. An example is displayed in 2(a). Here, each face contained within the triangular region defines both an $X$ and a $Z$ check. $Z$ and $X$ logical Pauli string operators [20] terminate at boundaries of the same color as displayed in the figure, or connect three boundaries of different colors. The code distance $d$ corresponds to the minimum weight of a non-trivial logical operator. In the example in 2(a) the distance is $d=5$.
The surface code [14, 15, 16] can be defined on a rectangular tiling with two different types of boundaries. The checks are placed in a checkerboard pattern. An example is shown in 2(b). Here, representatives of logical operators can be defined as strings that connect boundaries of the same type. In this example, horizontal strings correspond to $Z_{L}$ (connecting so-called rough boundaries) and vertical strings (connecting so-called smooth boundaries) to $X_{L}$. Therefore, the distance of the displayed example is $d=5$. We note that a rotated variant of the surface code, which is supported within a region enclosed by the dotted line in 2(b), leads to a smaller qubit overhead.
In its original formulation, the surface code does not admit transversal implementation of the logical $\mathrm{S}$ gates that are compatible with geometrically-local qubit connectivity. This issue can be circumvented by folding [23, 22] the surface code along the diagonal (orange nodes in 2(b) and 2(c)) such that faces of different check type are stacked on top of each other. Additionally, smooth and rough boundaries are also stacked, such that each boundary enables access to both logical $X_{L}$ and $Z_{L}$ representatives, similarly as for the color code.
@@ III-B Lattice Surgery
Quantum computation requires a universal gate set, such as $\{\mathrm{CNOT},\mathrm{H},\mathrm{T}\}$. In general, transversal implementation of the two-qubit $\mathrm{CNOT}{}$ gate requires a non-local qubit connectivity. Therefore, to implement $\mathrm{CNOT}{}$ gates with local connectivity, one may resort to lattice surgery [17].
@@ III-C Related Work
Let us conclude this section by reviewing how the resulting macroscopic compilation task is solved in related work. Surface code compilation with a universal or smaller gate set was explored in [7, 8, 9]. Furthermore, it has been shown that the complexity of optimizing a quantum circuit in the lattice surgery model is NP-hard [30]. Related work by Litinski [19] motivated further research [10, 11], though it was argued that commuting Clifford gates until the end of the circuit may limit the parallelism [12, 13]. Whether or not Clifford gates are commuted to the end may depend on the circuit being compiled [31]. The aforementioned works focus on surface codes only; consequently, fault-tolerant compilation for other codes remained underexplored, despite some early ideas [32].
Using color codes instead of surface codes offers several advantages. For instance, the color code has a higher qubit density, supports transversal Clifford gates and hosts both $X_{L}$ and $Z_{L}$ per boundary. This simplifies the routing to some extent by removing the constraint of connecting a specific choice of boundaries. However, decoding algorithms for the color code tend to perform worse [33, 34, 35, 36, 37]. Nevertheless, recent experiments with color codes [38, 39, 40, 41] and importantly, lattice surgery demonstration [42] underpin the potential of color codes as a viable fault-tolerant architecture.
Concerning the surface code, permitting a stacked planar architecture or relaxing the constraint on nearest-neighbor qubit connectivity enables to use the folded surface code [22, 23]. It retains many advantages of the color code while allowing straightforward decoding, the same way as for the standard surface code [15, 43]. However, to the best of our knowledge, no previous work on compilation with these codes exists.
@@ IV Formalism for Fault-Tolerant Compilation
On the physical level we assume that the local qubit connectivity is given by a graph (cf. Fig. 1a) that may have a constant number of layers of physical qubits such that qubits within each layer and between neighbouring layers can have geometrically local interactions. Note that such an architecture is less hardware-demanding than the one needed to implement bivariate bicycle codes [44, 45]; as such, it may be easier to realize with superconducting qubits [46, 47], as well as neutral atoms [48, 49] or trapped ions [50, 51]. Physical data and ancilla qubits (black and white in Fig. 1) are placed on the vertices and their connectivity is given by the edges. This is the basis for choosing the code substrate, which is a tiling $\mathcal{S}$ (Fig. 1b) that can be used to realize logical qubits encoded in some topological quantum code. The substrate allows us to view a concrete physical qubit connectivity abstractly. We use the substrate to illustrate the layout of data qubits and geometrically-local checks of the code (faces of $\mathcal{S}$).
@@
In this work we assume that the input circuit $\mathcal{C}$ is a logical quantum circuit that contains logical $\mathrm{CNOT}+\mathrm{T}$ gates. To obtain a universal gate set, single-qubit Clifford gates would be necessary as well, but they are not considered here since the computational hardness of the routing challenge for topological codes is dominated by multi-qubit operations [12, 13, 11, 8]. Moreover, our concrete choices of codes and substrates detailed below allow for transversal single-qubit Clifford operations.
[…]
The concrete design of the stabilizers of a snake requires special care as the protocol must be implemented in a fault-tolerant and distance-preserving way. We detail two concrete constructions for logical color code patches and color code snakes, and folded surface code patches and surface code snakes in Sec. VI-B and Sec. VI-C, respectively.
@@ V Macroscopic Compilation of $\mathrm{CNOT}+\mathrm{T}$ circuits
Figure 5: Considered layouts, i.e., a choice of $\mathcal{L}$ and $\mathcal{A}$ that determines the routing graph $\mathcal{R}$, and their asymptotic packing ratio $c$, which we define as the number of logical data qubit patches (assuming that the folded surface code forms one patch) to the total number of patches. Light blue and gray nodes depict logical data qubits and ancilla patches, respectively.
@@
In the following sections we detail how we can solve the routing and mapping problem for the macroscopic compilation, similar to [12, 13, 8, 10].
@@ V-A Shortest-First VDP Solving
Let us start by considering logical circuits with $\mathrm{CNOT}{}$ gates between disjoint pairs of qubits only. Given a routing graph $\mathcal{R}=(V_{\mathcal{R}},E_{\mathcal{R}})$ and a labeling of logical data patches $\Lambda^{*}$, one can formalize the search for paths as vertex disjoint path problem (VDP) [52, 53, 54]. The aim of the VDP is to find non-overlapping paths between pairs of vertices $C=\{(r_{1},r^{\prime}_{1}),...,(r_{k},r^{\prime}_{k})\mid r_{i}\neq r_{j},r^{\prime}_{i}\neq r^{\prime}_{j},r_{i}\neq r^{\prime}_{j}\,\forall i,j\}\subset V_{\mathcal{L}}\times V_{\mathcal{L}}$. The considered vertices in $V_{\mathcal{L}}$ are mapped to logical qubit labels via a choice of $\Lambda^{*}$. Hence, the pairs to connect—i.e., the elements of $C$—represent the control and target qubits of $\mathrm{CNOT}{}$ gates and the paths contain the snakes and logical ancillas to perform lattice surgery between pairs of logical data qubits. Overall, $C$ can be considered as quantum circuit consisting of $\mathrm{CNOT}{}$ gates.
@@
A simple greedy algorithm [52] to find solutions for the VDP is to start with an empty solution $W_{i}$ that is iteratively filled with paths $p_{j}$. We start by computing the shortest path with Dijkstra’s algorithm between each pair in $C$, choose the shortest among those paths and add it to the solution $W_{i}=W_{i}\cup\text{argmin}_{p_{i}\in\{\textsc{DIJKSTRA}(r_{j},r^{\prime}_{j})\mid(r_{j},r^{\prime}_{j})\in C\}}|p_{i}|$. Then the vertices in the chosen path $p_{i}$ are removed from $\mathcal{R}$. As pointed out in the beginning of the section, a substrate may require nontrivial constraints on paths on $\mathcal{R}$ to be valid such that Dijkstra may have to be adapted accordingly. This process of finding the shortest path and removing vertices is repeated until paths for all pairs in $C$ are found or no further non-overlapping path can be found. It is guaranteed that no overlapping paths can occur since consumed vertices are removed from $\mathcal{R}$. Running this shortest-first VDP subroutine yields a set of paths $W_{i}=\{p_{1},...p_{n}\}$ that correspond to logical two-qubit gates, which can be executed in parallel with lattice surgery on the substrate.
@@
In general, a $\mathrm{CNOT}{}$ quantum circuit contains multiple sets of pairs $C_{i}$ as we do not assume that all $\mathrm{CNOT}{}$ gates have disjoint qubit support. Thus a $\mathrm{CNOT}{}$ quantum circuit can be viewed as $\tilde{C}=\{C_{1},...,C_{m}\}$. Additionally, the routing graph $\mathcal{R}$ may not allow us to find a set of vertex-disjoint paths for a given $C_{i}$. Thus, the aforementioned subroutine must be embedded in an algorithm, which we call shortest-first VDP solving. We start by employing the shortest-first VDP subroutine for $C_{1}$ and add the resulting set of paths $W_{i}$ to the overall solution $C^{\prime}$. If not all pairs in $C_{1}$ can be connected by non-overlapping paths, the remaining elements from $C_{1}$ are copied to $C_{2}$. This may lead to $C_{2}$ containing gates with shared qubit labels. Thus, $C_{2}$ must be fixed by pushing problematic gates into $C_{3}$, such that $C_{2}$ contains gates with disjoint qubit labels only. Then, $C_{3}$ and the subsequent sets may be adapted in the same manner. Note that pushing requires us to respect the order of gates, as otherwise we could shuffle non-commuting gates. This shortest-first VDP subroutine together with the pushing is repeated until a path for each pair is found. The solution contains multiple sets of paths $C^{\prime}=\{W_{1},...W_{\Delta}\}$, where $\Delta$ is the depth of the final, compiled circuit.
@@
To include $\mathrm{T}$ gates in this picture, one can expand the shortest-first VDP subroutine to allow singleton tuples, i.e., single vertices $r_{\ell}\in V_{\mathcal{L}}$ corresponding to logical qubit labels via $\Lambda^{*}$. Additionally, the vertices of the magic state patches $V_{\mathcal{F}}$ must be taken into account in the routing problem formulation. We model magic state patches as logical patches that “receive” magic states after some fixed reset time from some adjacent magic state factory. The $\mathrm{T}$ gate can be implemented by gate injection from any available magic state patch. Thus, one can expand the shortest-first VDP subroutine by greedily computing all paths from the logical data qubit on which a $\mathrm{T}$ gate should be applied, $r_{\ell}$, to the magic state patches from $V_{\mathcal{F}}$, if their $\mathrm{T}$ state is available. The shortest paths between $r_{\ell}$ and the available magic state patches is compared to all other paths corresponding to elements in $C_{i}$ and the shortest is added to the solution as before. The availability of magic states is modeled by defining a reset time $t\in\mathbb{Z}$, which is initialized as positive integer and reduced by 111 after each completion of the shortest-first VDP subroutine. The factory is available when $t\leq 0$ and reset to the original value if the magic state was consumed. Note that is a pessimistic assumption as the units of time required to create a magic state may be less coarse grained.